# Jinhang Chai     Jianqing Fan

### v1

@@ Appendix C Proofs in Section 5
$\|\mathfrak{X}(W)\|_{\max}=\|X^{\top}W\|_{\max}\leq\|X^{\top}\|_{2,\infty}\cdot\|W\|_{2,\infty}=B\|W\|_{2,\infty}$
A direct application of 3.5 gives
@@
C.2 Proof of Theorem 5.3
The proof utilized Proposition C.1, which along with its proof, is given at the end of this section.
@@
Let $W=\widehat{\Sigma}-\Sigma$, we have from the Proposition C.1 that with probability at least $1-6p^{2}\exp(-ct^{2})$, $\|W\|_{\max}\leq\frac{t}{\sqrt{n}}$ and $\|W\|\leq p\|W\|_{\max}\leq\frac{pt}{\sqrt{n}}$.
@@
$\lesssim r\frac{p^{2}t^{2}}{n}+s\frac{t^{2}}{n}\lesssim\frac{rp^{2}t^{2}}{n}$
@@
$\|\widehat{\Sigma}_{L}-\Sigma_{L}\|^{2}_{F}+\|\widehat{\Sigma}_{S}-\Sigma_{S}\|^{2}_{F}=\mathcal{O}_{p}(\frac{rp^{2}\log^{2}p}{n})$
Proposition C.1.
@@
As $\mathbb{E}(X_{i}^{4})\leq C$, we have $\mathbb{E}|X_{i}|$ , $\mathbb{E}X_{i}^{2}X_{j}^{2}$, $\mathbb{E}X_{i}^{2}$ are all of constant order. Let $\max_{i,j}\{\mathbb{E}|X_{i}|,\mathbb{E}X_{i}^{2}X_{j}^{2},\mathbb{E}X_{i}^{2}\}\leq C_{0}$. Using the Winsorized concentration inequality established in Fan et al., (2016), we have when taking $\tau_{1}\asymp\sqrt{n}\gtrsim\sqrt{n\mathbb{E}X_{i}^{2}}$, it holds that
$\mathbb{P}(|\widehat{\mathbb{E}X_{i}}^{\top}-\mathbb{E}X_{i}|>t\frac{C_{0}}{\sqrt{n}})\leq 2\exp(-ct^{2}).$
Similarly for $X_{i}X_{j}$ we have when taking $\tau_{2}\asymp\sqrt{n}\gtrsim\sqrt{n\mathbb{E}X_{i}^{2}X_{j}^{2}}$
$\mathbb{P}(|\widehat{\mathbb{E}X_{i}X_{j}}^{\top}-\mathbb{E}X_{i}X_{j}|>t\frac{C_{0}}{\sqrt{n}})\leq 2\exp(-ct^{2})$
@@
$\leq$ $\mathbb{P}(|\mathbb{E}X_{j}|\cdot|\widehat{\mathbb{E}X_{i}}^{\top}-\mathbb{E}X_{i}|+|\widehat{\mathbb{E}X_{i}}^{\top}|\cdot|\widehat{\mathbb{E}X_{j}}^{\top}-\mathbb{E}X_{j}|+|\widehat{\mathbb{E}X_{i}X_{j}}^{\top}-\mathbb{E}X_{i}X_{j}|>t\frac{C_{0}}{\sqrt{n}})$
@@
By Weyl’s theorem and Davis-Kahan theorem, we have
$|\lambda_{l}-\widehat{\lambda}_{l}|\leq\|\widehat{\Sigma}_{L}-\Sigma_{L}\|$ (57)
and
@@
As the result is up to a rotational ambiguity, we can set the rotation matrix $H$ to be a diagonal matrix with $l^{th}$ diagonal element equal to $e_{l}$. It is straightforward to check that applying $H$ on $B$ helps align $B$ with $\widehat{B}$. Therefore without loss of generality, we can assume $e_{l}=1,\ \forall l$, and $H=\mathbf{I}$. It boils down to bound $\sum_{j=1}^{p}$
@@
Putting together we get
$\sum_{j=1}^{p}\|b_{j}-\widehat{b}_{j}\|^{2}$ $\lesssim r^{2}\frac{1}{p}\|\widehat{\Sigma}_{L}-\Sigma_{L}\|^{2}+rp\frac{\mu r}{p}\frac{1}{p}\|\widehat{\Sigma}_{L}-\Sigma_{L}\|^{2}$ (59)
$\lesssim r^{2}\frac{1}{p}\|\widehat{\Sigma}_{L}-\Sigma_{L}\|_{F}^{2}$ (60)
$\lesssim\frac{r^{3}p\log^{2}p}{n}$ (61)
Appendix D Extension to rectangular matrix estimand
@@
    $s\leq\bar{s}\leq\frac{\max\{p,q\}}{8c\bar{\mu}\bar{r}^{4}}$.
Corollary D.2.
Consider the case that $Y,L,S$ have dimension $p\times q$ instead of $p\times p$. Under Assumption 3.1,3.2,3.4, (4) still holds.
@@
Lemma D.3.
Let $P=U\Sigma V^{\top}$ and $Q=L\Lambda R^{\top}$ be the SVD of $P,Q\in\mathbb{R}^{p\times q}$, $U,L\in\mathcal{O}_{p,r},\ V,R\in\mathcal{O}_{q,r}$. Denote $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}\geq 0$ and $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{r}\geq 0$ as the diagonal elements of $\Sigma$ and $\Lambda$, respectively. Suppose $P$ and $Q$ both satisfy the $\mu$-incoherence condition 3.3, then we have for $\Delta=P-Q$,
@@
Without loss of generality, assume $p>q$. Define padding matrices $\tilde{P}=[P,0_{p,p-q}],\ \widetilde{Q}=[Q,0_{p,p-q}]$ and $\widetilde{\Delta}=\tilde{P}-\widetilde{Q}$. It is straightforward to verify that the SVD of $\tilde{P},\widetilde{Q}$ are $\tilde{P}=U\Sigma\widetilde{V}^{T}$ and $\widetilde{Q}=L\Sigma\widetilde{R}^{T}$, where $\widetilde{V}=\begin{bmatrix}V\\ 0_{p-q,r}\end{bmatrix}$ and $\widetilde{R}=\begin{bmatrix}R\\ 0_{p-q,r}\end{bmatrix}$.
By the definition of the incoherence, we have the incoherence of $\widetilde{V}$ and $V$ are equal and therefore, $\widetilde{V}\in\mathcal{O}_{p,r}$. Similarly, we have $\widetilde{R}\in\mathcal{O}_{p,r}$. Apply the key lemma B.1 with $P,Q$ replaced by $\tilde{P},\widetilde{Q}$, we get
@@
Appendix E Other Technical Lemmas
E.1 Proof of Lemma A.1
@@ Appendix E Other Technical Lemmas
E.2 Gilbert-Varshamov bound
Lemma E.1.

### v2

@@ Appendix E Proofs in Section 4
$\|\mathfrak{X}(W)\|_{\max}=\|X^{\top}W\|_{\max}\leq\|X^{\top}\|_{2,\infty}\cdot\|W\|_{2,\infty}=B\|W\|_{2,\infty}.$
A direct application of 2.5 gives
@@
E.2 Proof of Theorem 4.3
The proof utilized Proposition E.1, which along with its proof, is given at the end of this section.
@@
Let $W=\widehat{\Sigma}-\Sigma$, we have from the Proposition E.1 that with probability at least $1-6p^{2}\exp(-ct^{2})$, $\|W\|_{\max}\leq\frac{t}{\sqrt{n}}$ and $\|W\|\leq p\|W\|_{\max}\leq\frac{pt}{\sqrt{n}}$.
@@
$\lesssim r\frac{p^{2}t^{2}}{n}+s\frac{t^{2}}{n}\lesssim\frac{rp^{2}t^{2}}{n}.$
@@
$\|\widehat{\Sigma}_{L}-\Sigma_{L}\|^{2}_{F}+\|\widehat{\Sigma}_{S}-\Sigma_{S}\|^{2}_{F}=\mathcal{O}_{p}(\frac{rp^{2}\log^{2}p}{n}).$
Proposition E.1.
@@
As $\mathbb{E}(X_{1i}^{4})\leq C$, we have $\mathbb{E}|X_{1i}|$ , $\mathbb{E}X_{1i}^{2}X_{1j}^{2}$, $\mathbb{E}X_{1i}^{2}$ are all of constant order. Let $\max_{i,j}\{\mathbb{E}|X_{1i}|,\mathbb{E}X_{1i}^{2}X_{1j}^{2},\mathbb{E}X_{1i}^{2}\}\leq C_{0}$. Using the Winsorized concentration inequality established in Fan et al., (2016), we have when taking $\tau_{1}\asymp\sqrt{n}\gtrsim\sqrt{n\mathbb{E}X_{1i}^{2}}$, it holds that
$\mathbb{P}(|\widehat{\mathbb{E}X_{\cdot i}}-\mathbb{E}X_{1i}|>t\frac{C_{0}}{\sqrt{n}})\leq 2\exp(-ct^{2}).$
Similarly for $X_{1i}X_{1j}$, when taking $\tau_{2}\asymp\sqrt{n}\gtrsim\sqrt{n\mathbb{E}X_{1i}^{2}X_{1j}^{2}}$, it holds that
$\mathbb{P}(|\widehat{\mathbb{E}X_{\cdot i}X_{\cdot j}}-\mathbb{E}X_{1i}X_{1j}|>t\frac{C_{0}}{\sqrt{n}})\leq 2\exp(-ct^{2})$
@@
$\leq$ $\mathbb{P}(|\mathbb{E}X_{1j}|\cdot|\widehat{\mathbb{E}X_{\cdot i}}-\mathbb{E}X_{1i}|+|\widehat{\mathbb{E}X_{\cdot i}}|\cdot|\widehat{\mathbb{E}X_{\cdot j}}-\mathbb{E}X_{1j}|+|\widehat{\mathbb{E}X_{\cdot i}X_{\cdot j}}-\mathbb{E}X_{1i}X_{1j}|>t\frac{C_{0}}{\sqrt{n}})$
@@
By Weyl’s theorem, we have
$|\lambda_{l}-\widehat{\lambda}_{l}|\leq\|\widehat{\Sigma}_{L}-\Sigma_{L}\|.$ (59)
To prove (22), by the Davis-Kahan theorem (Davis & Kahan,, 1970) and the assumption on the minimum eigenvalue, it holds that there exists some $R\in{\cal O}_{r,r}$ such that
$\|VR-\widehat{V}\|\leq\frac{\sqrt{2}}{p}\|\widehat{\Sigma}_{L}-\Sigma_{L}\|$
Set $H=\Sigma^{-1/2}R\Sigma^{1/2}$. As $B=V\Sigma^{1/2}$ and $\widehat{B}=\widehat{V}\widehat{\Sigma}^{1/2}$, it follows that
$\|BH-\widehat{B}\|$ $=\|VR\Sigma^{1/2}-\widehat{V}\widehat{\Sigma}^{1/2}\|$
$\leq\|VR\Sigma^{1/2}-\widehat{V}\Sigma^{1/2}\|+\|\widehat{V}\Sigma^{1/2}-\widehat{V}\widehat{\Sigma}^{1/2}\|$
$\lesssim\sqrt{p}\|VR-\widehat{V}\|+\|\Sigma^{1/2}-\widehat{\Sigma}^{1/2}\|$
$\lesssim\frac{1}{\sqrt{p}}\|\widehat{\Sigma}_{L}-\Sigma_{L}\|+\frac{\|\widehat{\Sigma}_{L}-\Sigma_{L}\|}{\lambda_{r}^{1/2}}$
$\lesssim\frac{1}{\sqrt{\lambda_{r}}}\|\widehat{\Sigma}_{L}-\Sigma_{L}\|,$
where the last inequality follows from $\lambda_{r}\lesssim p$, as implied by the boundedness of of $\mathbb{E}X_{1i}^{2}$. Hence from Theorem 4.3 we have with probability at least $1-n^{-c}$ that
$\|BH-\widehat{B}\|\lesssim\frac{rp^{2}\log^{2}(p)\log n}{n\lambda_{r}},$
which completes the proof of (22).
To prove (23), we again invoke the Davis-Kahan theorem,
@@
As the result is up to a rotational ambiguity, we can set the rotation matrix $H$ to be a diagonal matrix with $l^{th}$ diagonal element equal to $e_{l}$. It is straightforward to check that applying $H$ on $B$ helps align $B$ with $\widehat{B}$. Therefore, without loss of generality, we can assume $e_{l}=1,\ \forall l$, and $H=\mathbf{I}$. It boils down to bound $\sum_{j=1}^{p}\|b_{j}-\widehat{b}_{j}\|^{2}$.
@@
Putting together we get with probability at least $1-n^{-c}$,
$\sum_{j=1}^{p}\|b_{j}-\widehat{b}_{j}\|^{2}$ $\lesssim r^{2}\frac{1}{p}\|\widehat{\Sigma}_{L}-\Sigma_{L}\|^{2}+rp\frac{\mu r}{p}\frac{1}{p}\|\widehat{\Sigma}_{L}-\Sigma_{L}\|^{2}$
$\lesssim r^{2}\frac{1}{p}\|\widehat{\Sigma}_{L}-\Sigma_{L}\|_{F}^{2}$
$\lesssim\frac{r^{3}p\log^{2}(p)\log n}{n}.$
Appendix F Extension to rectangular matrix estimand
@@
    $s\leq\bar{s}\leq\frac{\max\{p,q\}}{c\bar{\mu}\bar{r}^{4}}$.
Corollary F.2.
Consider the case that $Y,L,S$ have dimension $p\times q$ instead of $p\times p$. Under Assumption 2.1,2.2,2.4, (4) still holds.
@@
Lemma F.3.
Let $P=U\Sigma V^{\top}$ and $Q=L\Lambda R^{\top}$ be the SVD of $P,Q\in\mathbb{R}^{p\times q}$, $U,L\in\mathcal{O}_{p,r},\ V,R\in\mathcal{O}_{q,r}$. Denote $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}\geq 0$ and $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{r}\geq 0$ as the diagonal elements of $\Sigma$ and $\Lambda$, respectively. Suppose $P$ and $Q$ both satisfy the $\mu$-incoherence condition 2.3, then we have for $\Delta=P-Q$,
@@
Without loss of generality, assume $p>q$. Define padding matrices $\tilde{P}=[P,0_{p,p-q}],\ \widetilde{Q}=[Q,0_{p,p-q}]$ and $\widetilde{\Delta}=\tilde{P}-\widetilde{Q}$. It is straightforward to verify that the SVD of $\tilde{P},\widetilde{Q}$ are $\tilde{P}=U\Sigma\widetilde{V}^{\top}$ and $\widetilde{Q}=L\Sigma\widetilde{R}^{\top}$, where $\widetilde{V}=\begin{bmatrix}V\\ 0_{p-q,r}\end{bmatrix}$ and $\widetilde{R}=\begin{bmatrix}R\\ 0_{p-q,r}\end{bmatrix}$.
By the definition of the incoherence, we have the incoherence of $\widetilde{V}$ and $V$ are equal and therefore, $\widetilde{V}\in\mathcal{O}_{p,r}$. Similarly, we have $\widetilde{R}\in\mathcal{O}_{p,r}$. Apply the key lemma D.1 with $P,Q$ replaced by $\tilde{P},\widetilde{Q}$, we get
@@
Appendix G Other Technical Lemmas
G.1 Proof of Lemma C.1
@@ Appendix G Other Technical Lemmas
G.2 Gilbert-Varshamov bound
Lemma G.1.
@@ G.2 Gilbert-Varshamov bound
G.3 A simple lemma
The following is a simple lemma regarding relations between elements of $P$ and $\pi$, whose proof follows from simple algebra and is omitted.
Lemma G.2.
$P_{\min}\leq\pi_{\min}\leq\pi_{\max}\leq P_{\max},$